January 1 of the current year, Phobos Company acquired 10,000 shares of Investment in equity designated as a Fair Value through Other Comprehensive Income of Deimos Company at P 400,000 plus brokerage expenses of P 20,000. On March 1 of the current year, Deimos Company ordinary share was split on a 5 for 2 basis. On October 1, Deimos Company made a Special assessment of P 3.20 per share on all ordinary shareholder. Phobos Company accordingly paid the assessment. The fair value of December 31 amounted to P 30 per share:

Based on the above data answer the following:

1. The total number of shares at the end of the year

a. Nil c. 300,000

b. 140,000 d. 25,000

2. The unrealized gain to be presented in the other comprehensive income for the current year:

a. Nil c. 300,000

b. 140,000 d. 250,000




1
Expert's answer
2021-06-07T11:28:08-0400

Solution:

1.). Total shares outstanding before the share split is = 10,000 shares

Shares split ratio = 5:2, which means that for every 2 shares held by a shareholder, they will get 5 shares.

We first multiply the shares outstanding by five = 10,000 "\times" 5 = 50,000


Then we divide the number of shares by the two = "\frac{50,000}{2} = 25,000"


The total number of shares outstanding at the end of the year = 25,000 shares

 

2.). The unrealized gain to be presented in the other comprehensive income for the current year:

The unrealized gain = Fair value at the end of the year – Carrying amount

Fair value at the end of the year = Total shares outstanding x fair value per share

Fair value at the end of the year = 25,000 "\times" 30 = 750,000

Carrying amount = Acquisition value + brokerage fees + (assessment per share "\times" No. of outstanding shares)

Carrying amount = 400,000 + 20,000 + (3.20 "\times" 25,000)

Carrying amount = 420,000 + 80,000 = 500,000

The unrealized gain to be presented in the other comprehensive income for the current year = 750,000 – 500,000 = 250,000

 

The unrealized gain to be presented in the other comprehensive income for the current year = 250,000